### v1

# Diversity-Oriented Data Augmentation with Large Language Models

Abstract
Data augmentation is an essential technique in natural language processing (NLP) for enriching training datasets by generating diverse samples. This process is crucial for improving the robustness and generalization capabilities of NLP models. However, a significant challenge remains: Insufficient Attention to Sample Distribution Diversity. Most existing methods focus on increasing the sample numbers while neglecting the sample distribution diversity, which can lead to model overfitting. In response, we explore data augmentation’s impact on dataset diversity and propose a Diversity-oriented data Augmentation framework (DoAug). Specifically, we utilize a diversity-oriented fine-tuning approach to train an LLM as a diverse paraphraser, which is capable of augmenting textual datasets by generating diversified paraphrases. Then, we apply the LLM paraphraser to a selected coreset of highly informative samples and integrate the paraphrases with the original data to create a more diverse augmented dataset. Finally, we conduct extensive experiments on 12 real-world textual datasets. The results show that our fine-tuned LLM augmenter improves diversity while preserving label consistency, thereby enhancing the robustness and performance of downstream tasks. Specifically, it achieves an average performance gain of $10.52\%$, surpassing the runner-up baseline with more than three percentage points.
Diversity-Oriented Data Augmentation with Large Language Models
Zaitian Wang1,2, Jinghan Zhang3, Xinhao Zhang3, Kunpeng Liu3, Pengfei Wang1,2, Yuanchun Zhou1,2 1Computer Network Information Center, CAS, 2University of Chinese Academy of Sciences, 3Portland State University wangzaitian23@mails.ucas.ac.cn
1 Introduction
AI methods have demonstrated immense capabilities, often surpassing human abilities and traditional techniques across various natural language processing (NLP) tasks. This success largely hinges on the availability of high-quality datasets, which enable AI models to uncover intrinsic patterns and drive their effectiveness in real-world applications. However, training on inferior datasets can significantly degrade model performance, particularly when applied to test data or real-world scenarios Wang et al. (2024b). As AI technology advances, especially with large language models (LLMs), the demand for high-quality datasets has become more pronounced. To effectively train NLP models, a high-quality dataset should be (1) Large: a sufficient number of samples is crucial to reflect the diversity and complexity of human language. Large datasets help prevent overfitting, ensuring the trained AI model generalizes well to unseen data. With more data points, the model can learn various patterns and relationships, which enhances its robustness and reliability; (2) Coherent: the mapping between data and labels must be accurate and consistent. Coherent datasets ensure that each data point is correctly labeled, providing the model with reliable information for learning. Incoherent datasets, with mislabeled or inconsistent data, can confuse the model and degrade its performance. Consistency in labeling also aids in the reproducibility of results and the interpretability of the model’s predictions; (3) Diverse: a diverse dataset ensures NLP models learn a broad spectrum of linguistic patterns, enhancing robustness across real-world conditions. This includes variations such as dialects, tones, formality, or domain-specific terms. Exposure to such diversity helps models generalize better, avoiding over-reliance on narrow language subsets. Additionally, it improves adaptability to unexpected inputs while reducing biases tied to certain language styles.
Data augmentation is an efficient technique for increasing the number of training samples by modifying existing dataset samples Wang et al. (2024b). It allows for the rapid generation of large-scale datasets without the need for additional data collection and has been successfully applied in various domains, including textual data. Data augmentation for textual data often changes the wording or reshapes the structure of a sentence. Many early and simple methods achieve this by randomly perturbing the textual samples at the word level Wei and Zou (2019); Karimi et al. (2021). While certain operations prove effective, some introduce noise that compromises label integrity (e.g. deleting a “not”) or generates almost redundant samples, ultimately failing to improve dataset quality or promote diversity. Recent advances in LLMs have demonstrated unprecedented power in text understanding and generation Radford et al. (2018). The capacity of these state-of-the-art (SOTA) generative language models establishes a new and promising paradigm of textual data augmentation Anaby-Tavor et al. (2020). Generative models enable large-scale acquisition of textual data while preserving the coherence of augmented datasets by generating texts with similar meanings to the original sentences Dai et al. (2025). However, most existing generative methods focus primarily on enlarging dataset size, with limited consideration of how the augmentation process affects dataset diversity. Adequate attention to maintaining and enhancing dataset diversity is vital for developing AI-ready, high-quality datasets, making it a crucial challenge in textual data augmentation.
Along this line, we propose a Diversity-oriented data Augmentation approach (DoAug) using an LLM to paraphrase sentences. The LLM is first fine-tuned on a paraphrase dataset and taught to rewrite sentences. By instructing the LLM to function as a paraphraser, we can use it to alter sentence expressions while preserving their essential meaning. In this way, we ensure the affinity between the original and augmented samples, minimizing the influence of data augmentation on dataset coherence. To enhance the dataset diversity through data augmentation, we further proposed a diversity-oriented fine-tuning method. We construct a preference dataset that chooses the more diverse paraphrases while rejecting repetitive ones. Then the LLM paraphraser is fine-tuned on the preference dataset with the DPO algorithm Rafailov et al. (2024) to encourage greater diversity in its generated samples. We also adopt a coreset selection method to focus on only the most important samples from the dataset to reduce the computational overhead and costs of running LLMs. Finally, we conduct extensive experiments on 12 textual benchmark datasets to verify the effectiveness of DoAug. Experimental results show that our proposed method can remarkably enhance the diversity of the augmented dataset on 5 measurements while maintaining high affinity compared with the original dataset. We also investigate the implications of this increased diversity on model performance in downstream tasks and observe significant improvements in model performance when trained on the augmented datasets.
In summary, the contributions are as follows:
(1) We propose a data augmentation framework, DoAug, that incorporates and explicitly encourages diversity, an important yet often neglected factor in high-quality datasets;
(2) The framework trains and employs an LLM as a paraphraser to generate synthetic data with high affinity, ensuring the coherence of the augmented datasets;
(3) We introduce a diversity-oriented fine-tuning method that trains the LLM augmenter on a preference dataset with the DPO algorithm the boost generation diversity of the LLM;
(4) Extensive experiments conducted on 12 datasets demonstrate that DoAug significantly benefits learning performance by increasing dataset diversity while maintaining coherence.
2 Related Work
2.1 Textual Data Augmentation
Textual data augmentation revolves around perturbing the wording and syntax of existing sentences to create more modified samples. Some early and simple methods propose to randomly replace, remove, insert, and swap characters or words at certain ratios in a sentence Belinkov and Bisk (2018); Wei and Zou (2019). Some more sophisticated methods modify sentences by using alternative syntax Min et al. (2020). Language models can in turn act as effective tools for textual data augmentation. For example, Back-translation Sennrich (2015) first translates sentences from the source language (e.g. English) to an intermediary (e.g. Chinese) and then translates the intermediary sentence back to the source language. Substitute Word using BERT Kumar et al. (2020) masks certain words in the original sentences and uses a BERT model to predict masked words. They utilize the subtle differences made by the translator or the unmasking process and perturb the wording or syntax while keeping the meanings untouched. The recent emergence of LLM has given birth to a series of new approaches Anaby-Tavor et al. (2020); Cai et al. (2023); Ding et al. (2024); Wang et al. (2024a). AugGPT Dai et al. (2025), for example, prompts the state-of-the-art ChatGPT model to rewrite sentences in the dataset and preserves dataset coherence after data augmentation. Self-LLMDA Li et al. (2024) automatically generates and selects the most suitable instruction to prompt the LLM to generate augmented samples. However, these aforementioned methods neglect the impact on dataset diversity, failing to ensure the diversity trait of producing high-quality datasets. The effect of LLM augmentation diversity is discussed in Cegin et al. (2023, 2024), where three types of prompt-based diversity incentives are proposed.
2.2 Dataset Diversity Evaluation
The evaluation of dataset diversity is increasingly popular as the size of available training data stunningly explodes, which makes it vital to maintain a minimized redundancy in the dataset to avoid repetitive training, saving the cost and time consumption and avoiding overfitting. Though its definition is not yet unified, many metrics are used across research. Tevet and Berant (2021) systematically studies the evaluation of text data diversity, which includes token-level metrics, embedding-level metrics, and human evaluations. Lai et al. (2020) proposes three dataset diversity metrics in the embedding space and investigates how these metrics change in different text datasets. Yu et al. (2022) proposes another three diversity metrics and discusses how improving dataset diversity helps enhance learning generalization, even when the total size of the dataset is reduced. Gontijo-Lopes et al. (2020) jointly investigates the role of data diversity and affinity in data augmentation, demonstrating that model performance benefits from improvements in both measures. Diversity has been considered in the design of several data augmentation methods Malandrakis et al. (2019); Liu et al. (2021), however, it has not yet been integrated with coherence-ensured and LLM-based data augmentation methods such as AugGPT Dai et al. (2025).
Figure 1: An overall framework of DoAug.
3 Methodology
3.1 Problem Formulation
Given a parameterized data augmenter $f_{\theta}$, the data augmentation process is expressed as $f_{\theta}:\mathcal{S}=\{\mathbf{X},\mathbf{t}\}\rightarrow\mathcal{\tilde{S}}=\{\mathbf{\tilde{X}},\mathbf{\tilde{t}}\}$, where $\mathcal{S}$ is the original dataset composed of the feature vectors $\mathbf{X}$ and target labels $\mathbf{t}$, and $\mathcal{\tilde{S}}$ is the augmented dataset Wang et al. (2024b). For a diversity metric $D$, the diversity values of the original and the augmented datasets are $D(\mathcal{S})$ and $D(\mathcal{\tilde{S}})$, respectively, and the diversity gain of that augmentation is defined as $\Delta D(\mathcal{S};\theta)=D(\mathcal{\tilde{S}})-D(\mathcal{S})$.
DoAug aims to optimize the parameter $\theta^{\ast}$ for the data augmenter to maximize the diversity gain after augmentation:
$\vspace{-3mm}\theta^{\ast}=\mathop{\arg\max}_{\theta}\mathbbm{E}\Delta D(\mathcal{S};\theta)$ (1)
When training models on the original and augmented datasets, their respective performances are evaluated by a performance metric $P$, resulting $P(\mathcal{S})$ for the original dataset and $P(\mathcal{\tilde{S}})$ for the augmented dataset. The performance gain is defined as $\Delta P(\mathcal{S};\theta)=P(\mathcal{\tilde{S}})-P(\mathcal{S})$. Moreover, by optimizing and employing the augmenter with maximum diversity gain, DoAug expects to achieve maximum performance gain under fixed conditions:
$\vspace{-3mm}\Delta P(\mathcal{S};\theta^{\ast})=\mathop{\max}_{\theta}\Delta P(\mathcal{S};\theta)$ (2)
3.2 Framework Overview
DoAug trains and employs an LLM capable of generating diverse paraphrases for data augmentation to enlarge the size of textual datasets, maintain coherence, and enhance diversity. As Figure 1 shows, the framework is organized as follows:
$\bullet$ An LLM is first trained on a paraphrase dataset through supervised instruction fine-tuning, enabling it to function as a paraphraser that rewrites sentences while preserving the original semantics;
$\bullet$ The LLM paraphraser is then trained on a constructed preference dataset with the DPO (Direct Preference Optimization) algorithm that encourages diverse generation samples;
$\bullet$ For a given textual dataset, a coreset of samples is selected based on their importance, serving as the source sentences for paraphrase augmentation;
$\bullet$ Generated paraphrases are ranked and sampled according to their diversity, with the most diverse paraphrases integrated into the augmented dataset with original coreset samples.
3.3 Diverse Paraphraser Fine-tuning
The proposed data augmentation method is constructed upon a general-purpose LLM. Pre-trained on a vast amount of corpus, LLMs now boast human-level understanding and generation abilities Ouyang et al. (2022); Dubey et al. (2024). We leverage these abilities of an LLM and use it as a tool for our data augmentation process. The LLM is fine-tuned as a paraphraser that can rewrite sentences with alternative expressions while maintaining the original semantics. The LLM paraphraser is further fine-tuned to produce more diverse generation results and cover more linguistic alternatives. On the one hand, the change in sentence expressions can introduce diversity to the dataset. On the other hand, since the LLM is capable of capturing the semantics of the sentence and is prompted only to paraphrase the sentence, the coherence of the augmented data-label mapping is preserved.
3.3.1 LLM Paraphraser Training with PEFT
To fully leverage the understanding and generation abilities of the LLM, we use supervised fine-tuning (SFT) to train it to follow instructions to paraphrase existing sentences. Since the SFT phase of LLM training is heavily computation-consuming, we use the Parameter-Efficient Fine-Tuning (PEFT) technique to reduce the size of trainable parameters updated in the back-propagation pass to save the computation cost. Specifically, we adopt the Low-Rank Adaptation (LoRA) approach Hu et al. (2021). Given a pre-trained LLM with weights $W_{0}\in\mathbb{R}^{d\times k}$, LoRA represents its update $\Delta W$ with $BA$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, and the rank $r\ll min(d,k)$. During training, $W_{0}$ is frozen and excluded from the gradient update, while $A$ and $B$ are updated instead. After training, $W_{0}$ and $\Delta W=BA$ are multiplied with the same input, and their outputs are summed, as in:
$h=W_{0}x+\Delta Wx=W_{0}x+BAx\vspace{-0.15cm}$ (3)
In the SFT phase, we sample a subset $\mathcal{D}_{\text{SFT}}$ from ChatGPT Paraphrases dataset [^1] https://huggingface.co/datasets/humarin/chatgpt-paraphrases to train the LLM.
3.3.2 LLM Generation Diversity Enhancement with DPO
To align LLM generations to human preferences, recent LLMs adopt RLHF in their training process, which involves the PPO algorithm and a reward model Ouyang et al. (2022); Zhang et al. (2024). However, fitting a reward model brings extra computation costs, and the gap between its prediction and actual human preference also poses threats to the effect of PPO. As an alternative, the DPO algorithm directly optimizes the LLMs’ generation policy without training an additional reward model Rafailov et al. (2024).
Preference Dataset Construction.
To construct a dataset $\mathcal{D}_{\text{DPO}}$ for DPO training, we sample another subset from the original paraphrase dataset. Each original sentence corresponds to 5 paraphrases, and we embed the original sentence $x$ and paraphrased sentences $[y_{1},...,y_{5}]$ and calculate the Euclidean distances in the embedding space $\mathcal{E}$:
$dist(y_{i},x)=\sqrt{(e_{y_{i}}-e_{x})^{2}},\vspace{-0.25cm}$ (4)
where $e_{x}=\mathcal{E}(x)$. The paraphrase with maximum distances is considered the most diverse among possible generation results and used as the “chosen” (preferred) output. In contrast, the most similar is taken as the least varied generation and used as the “rejected” (dispreferred) one. This preference construction process is formulated in Eq. 5:
$y_{w}=\arg\max_{y_{i}}dist(y_{i},x)$ (5)
$y_{l}=\arg\min_{y_{i}}dist(y_{i},x),$
where $y_{w}$ is the preferred paraphrase out of the pair $(y_{w},y_{l})$.
Training Objective.
The goal of DPO training is to maximize the probability of generating the preferred output and minimize the probability of generating the dispreferred output. Unlike the PPO algorithm which requires a reward model and a reinforcement learning phase, DPO derives its objective by solving the optimal solution of PPO’s optimization problem, as shown in Eq. 6:
$\vspace{-3mm}\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{\text{DPO}}}\bigg{[}\log\sigma\bigg{(}\\ \beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\left.\left.-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right],$ (6)
where $\pi_{\theta}$ denotes the generation probability of the current LLM, $\pi_{\text{ref}}$ denotes that of the SFT model, and hyper-parameter $\beta$ controls the deviation from the SFT model.
3.3.3 Diversity-oriented Sampling
For each input sentence, we use beam search to generate $K$ sequences from the LLM’s output logits. We then rank these sequences based on their distances from the original sentences according to Eq. 4. Only the most distant sentences are retained to reduce redundancy and prevent overfitting.
Algorithm 1 Diversity-oriented and Coreset-focused Data Augmentation
0:  An LLM $f_{\theta}$, a paraphrase dataset $\mathcal{D}_{\text{SFT}}$, a preference dataset $\mathcal{D}_{\text{DPO}}$, and a target textual dataset $\mathcal{S}$
1:  Train the original LLM $f_{\theta}$ on the paraphrase dataset $\mathcal{D}_{\text{SFT}}$ with SFT
2:  Fine-tune the LLM on the preference dataset $\mathcal{D}_{\text{DPO}}$ by optimizing the loss function in Eq. 6
3:  Initialize empty $\mathcal{S}^{\prime}$
4:  for all $(x,t)\in\mathcal{S}$  do
5:     Calculate importance score $s$ for $x$
6:     $\mathcal{S}^{\prime}$.append($(x,t,s)$)
7:  end for
8:  Rank samples in $\mathcal{S}^{\prime}$ according to the score key $s$
9:  Split $\mathcal{S}^{\prime}$ with ratio $r_{\text{augment}}:r_{\text{retain}}:r_{\text{prune}}$ to $\mathcal{S}_{\text{augment}}$, $\mathcal{S}_{\text{retain}}$, and $\mathcal{S}_{\text{prune}}$
10:  Initialize empty $\mathcal{\tilde{S}}$
11:  for all $(x,t,s)\in\mathcal{S}_{\text{augment}}$ do
12:     $y=f_{\theta}(x)$
13:     $\mathcal{\tilde{S}}$.append($(x,t)$)
14:     $\mathcal{\tilde{S}}$.append($(y,t)$)
15:  end for
16:  $\mathcal{\tilde{S}}=\mathcal{\tilde{S}}\cup\mathcal{S}_{\text{select}}$
17:  return  Augmented dataset $\mathcal{\tilde{S}}$
3.4 Selective Coreset Data Augmentation
Given the time consumption and computation cost of LLM-based data augmentation methods, it is non-trivial to recognize the most important samples and constrain the target for data augmentation to these samples. First, we train the downstream task model on the dataset and collect training dynamics and post-training post-training metrics. Then we calculate the EL2N Paul et al. (2021), entropy Coleman et al. (2019), variance Swayamdipta et al. (2020), and AUM Pleiss et al. (2020) score to evaluate sample importance. We use score monotonic selection and coverage-centric selection (CCS) Zheng et al. (2022) to derive the coresets. DoAug performs a hierarchical coreset selection to prune some low-importance samples, retain the middle-importance samples, and augment the high-importance samples. Only samples of high importance are used as the seeds for data augmentation. The original sentences in the high-importance coreset and their paraphrases are combined with the middle-importance samples, composting the final results of our data augmentation process, as presented in Algorithm 1.
4 Experiments
4.1 Experiment Settings
Evaluation Criterion. We evaluate diversity and affinity for data distribution while measuring performance on downstream tasks for effectiveness.
Diversity. To comprehensively evaluate the effect of the proposed model DoAug on dataset diversity, we adopt several measures to evaluate the augmented dataset’s diversity:
$\bullet$ Distance: assesses the average distance between samples as follows: ${Distance(\mathcal{S})=\frac{1}{|\mathcal{S}|}\sum_{x_{i},x_{j}\in\mathcal{S}}\sqrt{(e_{x_{i}}-e_{x_{j}})^{2}}}$, where $e_{x}=\mathcal{E}(x)$ is the embedding of sample $x$ in the embedding space $\mathcal{E}$, and a larger distance indicates greater diversity.
[…]
Performance on downstream task. Following the practice in existing research on textual data augmentation, we train a $\text{BERT}_{\text{base}}$ model Kenton and Toutanova (2019) with a classification head on the original and augmented datasets to evaluate the effect of our proposed data augmentation approach. We report the prediction accuracy scores on each dataset to measure downstream task performance.
Datasets. We conduct extensive experiments on 12 NLP datasets to verify the effectiveness of DoAug. As specified in Appendix A, our selection of datasets covers a wide range of text classification tasks and includes datasets with different characteristics. Following the settings in Yoo et al. (2021), we sample a subset (1.2K samples) from the full dataset to unify the evaluation settings, enable a fair comparison between methods, and simulate a low-resource condition where data augmentation is of significant necessity.
Baseline Methods. We compare DoAug with eleven representative data augmentation methods. (1) OCR and (2) Keyboard perform common OCR or typing errors at the character level. (3) EDA randomly inserts, deletes, replaces, or swaps words in the sentences to create more samples. (4) Back-translation (BT) involves translating the source sentences to an intermediary language. (5) Unmask randomly replaces words with [MASK] and predicts the masked words with the BERT model. (6) AugGPT directly prompts ChatGPT for paraphrases without further fine-tuning Dai et al. (2025). (7) Grammar and (8) Spelling are two exemplar methods selected by Self-LLMDA Li et al. (2024), which prompt the LLM to simulate common grammatical variation or spelling errors made by humans. (9) Chain, (10) Hint, and (11) Taboo generate paraphrases with three different diversity incentives Cegin et al. (2024).
ANLI ChemProt CoLA MNLI MPQA MRPC RCT RTE SST-2 SUBJ Symptoms Yelp Avg. Gain
Original $35.75_{1.68}$ $58.33_{4.67}$ $74.56_{1.11}$ $42.81_{2.33}$ $89.17_{0.47}$ $76.50_{3.09}$ $71.62_{2.49}$ $53.61_{2.45}$ $86.97_{1.00}$ $\underline{95.75}_{0.23}$ $74.06_{9.86}$ $51.48_{6.81}$
OCR $34.43_{1.50}$ $63.84_{3.14}$ $73.12_{1.82}$ $53.94_{8.10}$ $88.83_{0.48}$ $75.49_{4.48}$ $\underline{79.66}_{1.02}$ $56.06_{3.92}$ $86.91_{0.74}$ $95.25_{0.25}$ $86.12_{2.72}$ $55.46_{1.04}$ $4.73\%$
Keyboard $34.99_{1.53}$ $65.78_{2.28}$ $72.69_{1.19}$ $52.88_{4.44}$ $88.98_{0.43}$ $77.99_{2.74}$ $79.42_{0.72}$ $57.58_{4.23}$ $86.12_{0.66}$ $95.25_{0.37}$ $87.24_{1.09}$ $55.44_{0.62}$ $5.40\%$
EDA $34.91_{2.52}$ $64.24_{2.45}$ $73.07_{0.95}$ $55.90_{3.21}$ $89.12_{0.60}$ $79.44_{2.11}$ $77.17_{1.28}$ $55.96_{3.85}$ $87.48_{1.07}$ $95.63_{0.27}$ $89.11_{1.45}$ $55.15_{1.34}$ $6.68\%$
BT $34.39_{0.78}$ $67.72_{2.83}$ $70.42_{2.03}$ ${56.40}_{3.19}$ $\underline{89.73}_{0.38}$ $78.55_{2.62}$ $76.36_{1.09}$ $53.52_{3.67}$ $85.89_{1.42}$ $95.16_{0.39}$ $89.43_{0.59}$ ${55.56}_{0.85}$ $6.26\%$
Unmask $36.04_{1.19}$ $69.36_{1.85}$ $74.11_{1.10}$ $54.60_{5.77}$ $88.87_{0.45}$ $\underline{80.15}_{1.27}$ $79.01_{0.87}$ $55.85_{3.97}$ $87.08_{1.00}$ $95.22_{0.21}$ $89.92_{0.55}$ $55.10_{0.69}$ $6.59\%$
AugGPT $\underline{36.43}_{1.04}$ $65.73_{2.80}$ $\underline{75.17}_{1.35}$ $53.77_{2.61}$ $89.67_{0.34}$ $75.25_{2.12}$ ${78.90}_{0.70}$ $54.87_{2.51}$ $\underline{87.63}_{0.60}$ $95.44_{0.31}$ $79.25_{3.33}$ $55.47_{0.64}$ $5.64\%$
Grammar $35.39_{1.55}$ $68.64_{2.47}$ $72.16_{1.45}$ $56.53_{1.30}$ $89.46_{0.42}$ $78.90_{1.83}$ $77.35_{0.68}$ $54.40_{3.58}$ $86.62_{1.56}$ $94.98_{0.25}$ $\underline{89.98}_{0.43}$ $55.27_{0.82}$ $5.88\%$
Spelling $35.98_{1.35}$ $68.69_{2.06}$ $72.09_{1.66}$ $56.76_{1.04}$ $88.96_{0.52}$ $79.12_{1.77}$ $78.95_{0.66}$ $57.40_{3.09}$ $86.42_{0.89}$ $95.20_{0.42}$ $89.48_{0.92}$ $54.76_{0.99}$ $6.49\%$
Chain $35.25_{1.82}$ $68.75_{2.66}$ $72.32_{0.86}$ $56.23_{1.64}$ $89.10_{0.41}$ $75.88_{2.85}$ $79.31_{0.54}$ $54.98_{5.42}$ $86.79_{0.82}$ $95.14_{0.36}$ $89.24_{0.55}$ $55.50_{0.66}$ $5.77\%$
Hint $36.19_{1.24}$ $68.67_{1.96}$ $72.25_{1.47}$ $56.69_{1.47}$ $89.13_{0.67}$ $78.75_{2.07}$ $78.44_{0.41}$ $55.78_{2.98}$ $86.80_{0.74}$ $95.00_{0.31}$ $89.58_{1.35}$ $55.88_{0.32}$ $6.43\%$
Taboo $35.83_{1.75}$ $\underline{69.66}_{1.79}$ $72.90_{1.21}$ $\underline{57.26}_{0.93}$ $89.34_{0.29}$ $76.74_{2.30}$ $78.48_{0.41}$ $\mathbf{58.01}_{3.41}$ $86.74_{1.43}$ $95.12_{0.34}$ $89.40_{0.64}$ $\underline{56.30}_{0.71}$ $\underline{6.76\%}$
DoAug $\mathbf{38.46}_{2.51}$ $\mathbf{70.22}_{1.76}$ $\mathbf{75.62}_{0.51}$ $\mathbf{59.76}_{1.02}$ $\mathbf{89.78}_{0.29}$ $\mathbf{80.97}_{3.45}$ $\mathbf{80.10}_{0.63}$ $\underline{56.05}_{2.72}$ $\mathbf{88.64}_{0.80}$ $\mathbf{95.80}_{0.19}$ $\mathbf{90.74}_{1.85}$ $\mathbf{56.57}_{0.49}$ $\mathbf{10.52\%}$
Table 1: Prediction accuracy of models trained on augmented datasets. The best results are highlighted with the bold font, and runner-ups are underlined. We report the mean performance and standard deviation and the results are averaged on ten random seeds.
Figure 2: Diversity, affinity, and performance achieved by DoAug and baseline methods. Results are averaged on 12 datasets and the diversity rankings are further averaged on 5 metrics in this diagram. Smaller number for the rankings indicates better results.
4.2 Overall Results
To verify the effectiveness of DoAug, we evaluate downstream task accuracy alongside the diversity and affinity of the augmented dataset. We report the rankings of performance, diversity, and affinity averaged on 12 datasets achieved by DoAug and 11 baseline methods in Figure 2. From these results, we have the following observations: (1) DoAug achieves the highest performance on downstream tasks compared to other SOTA data augmentation methods, as indicated by the color bar in Figure 2. This demonstrates the high quality and superior adaptability of the datasets generated by our proposed method in real-world applications. (2) DoAug achieves the highest diversity score and outperforms all other baseline methods. This implies that DoAug effectively improves dataset diversity. (3) DoAug achieves a considerably high position on the affinity rankings, indicating that the sample semantics are preserved to the greatest extent possible. In sum, DoAug achieves the top position in the combined dataset diversity and affinity rankings. Additionally, it achieves the best downstream task performance, indicated by the lightest yellow.
Distance $\uparrow$ Dispersion $\uparrow$ Radius $\uparrow$ Homogeneity $\uparrow$ Vocabulary $\uparrow$ Average Rankings $\downarrow$
Original 8.9458 (11) .1911 (13) .5352 (6) .8834 (11) 4737 (8) 9.8
OCR 8.8516 (13) .1963 (9) .5198 (13) .8947 (9) 5764 (2) 9.2
Keyboard 8.9439 (12) .1978 (7) .5233 (12) .8990 (8) 5939 (1) 8.0
EDA 9.1927 (6) .2090 (3) .5247 (11) .9010 (7) 4416 (11) 7.6
BT 9.2761 (4) .2050 (4) .5345 (7) .8744 (12) 4782 (7) 6.8
Unmask 9.0021 (10) .1932 (12) .5344 (8) .9041 (6) 4158 (13) 9.8
AugGPT 9.1522 (8) .2005 (6) .5340 (9) .9094 (3) 4585 (10) 7.2
Grammar 9.5250 (2) .2121 (2) .5384 (3) .8332 (13) 4316 (12) 6.4
Spelling 9.0346 (9) .1951 (11) .5315 (10) .9080 (4) 4917 (5) 7.8
Chain 9.1946 (5) .1964 (8) .5385 (2) .9099 (2) 4974 (4) 4.2
Hint 9.3608 (3) .2037 (5) .5388 (1) .8946 (10) 4796 (6) 5.0
Taboo 9.1566 (7) .1953 (10) .5379 (4) .9148 (1) 4690 (9) 6.2
DoAug 9.5308 (1) .2154 (1) .5355 (5) .9054 (5) 5321 (3) 3.0
Table 2: 5 Diversity metrics averaged on 12 datasets.
4.3 Performance, Diversity, and Affinity
4.3.1 Performance Gains
The full results for BERT classification performance on original and augmented datasets are presented in Table 1. The results show that DoAug surpasses all baseline methods on average. Specifically, it outperforms the baseline methods on 11 out of 12 datasets except on RTE. DoAug achieves performance gain of $10.52\%$ on average, surpassing the runner-up method, namely Taboo, with 3.76 percentage points.
4.3.2 Diversity Gain
We demonstrate the diversity gains in terms of all 5 diversity metrics along with their rankings achieved by DoAug and baseline methods in Figure 2. DoAug ranks the top on the chart with an average ranking of 3.0. Specifically, it achieves the best for the Distance and Dispersion metrics. For the Vocabulary metric, OCR and Keyboard include large amounts of out-of-vocabulary tokens due to character level perturbation. The three baselines with diversity incentives, namely Chian, Hint, and Taboo, also achieve reasonably good diversity gain, in line with the results of Cegin et al. (2024).
Figure 3: Affinity scores of DoAug and 10 baseline methods. The scores are averaged on 12 datasets.
4.3.3 Affinity and Paraphrase Validity
We present the affinity of DoAug and baseline methods in Figure 3, where DoAug outperforms other methods except Unmask, whose affinity score is 1.95. The Unmask method generates augmentation by replacing randomly selected words with “[MASK]” and predicts the masked words with the BERT model. Since the augmented samples are recovered from the BERT embeddings of the corrupted original samples and we use the BERT embeddings to calculate affinity, it is reasonable to yield extremely high affinity scores. Following Cegin et al. (2023, 2024), we also investigate paraphrase validity at the sample level. We prompt the DeepSeek-V3 model to check if the paraphrases are semantically similar to the original samples and adhere to the original labels. Results show that $97\%$ paraphrases are valid, suggesting the data augmenter introduces negligible noises to the dataset.
4.4 Ablation Studies
To verify the effectiveness of our proposed method, we conduct ablation studies to show that all components in the method framework contribute to the final performance gains, as shown in Table 3, where w/o Coreset refers to applying augmentation on a random subset of the dataset without deriving a coreset of importance samples, w/o Selective refers to augmenting samples in both $\mathcal{S}_{\text{retain}}$ and $\mathcal{S}_{\text{augment}}$ instead of only augmenting the latter, w/o Aug refers to using the coreset directly for training without data augmentation, w/o DPO refers to using the LLM paraphraser from the SFT stage for data augmentation, w/o DS refers to removing the diversity-based sampling module, and w/o DPO|DS removes both steps. Thus, all components in our method framework contribute to the final performance gains. As Figure 4 shows, we also study the effect of diversity-oriented fine-tuning and diversity-based sampling on the dataset diversity, demonstrating that all proposed components are effective.
ANLI Ch.Pr. CoLA MNLI MPQA MRPC RCT RTE SST-2 SUBJ Sympt. Yelp
w/o Coreset 35.32 64.24 71.94 54.51 89.32 73.10 77.23 53.52 87.42 95.19 87.58 55.87
w/o Selective 35.94 66.40 73.35 52.14 89.14 75.46 77.23 55.69 87.90 95.49 89.95 56.27
w/o Aug 37.82 64.86 72.82 43.89 89.43 78.65 76.74 55.69 86.75 95.70 86.06 53.74
w/o DPO|DS 37.64 70.18 75.52 59.04 89.72 79.90 78.42 54.87 87.99 95.47 90.08 56.50
w/o DPO 37.32 69.49 74.80 58.75 89.34 80.67 78.43 55.37 88.15 95.37 90.66 56.52
w/o DS 36.85 69.62 75.49 59.32 89.53 80.18 78.56 53.70 88.14 95.51 90.65 56.36
DoAug 38.46 70.22 75.62 59.76 89.78 80.97 80.10 56.05 88.64 95.80 90.74 59.67
Table 3: Ablation study on model performance gains.
Figure 4: Ablation study on diversity gains
GPT-2 (137M) T5-large (738M)
CoLA MNLI RCT CoLA MNLI RCT
Original 67.64 50.03 76.53 78.94 50.06 78.98
OCR 65.51 54.42 77.70 77.79 56.35 81.77
Keyboard 65.23 54.51 77.37 77.06 56.40 81.65
EDA 66.60 54.13 76.89 78.59 51.58 80.79
BT 63.43 53.89 77.10 76.22 54.70 80.80
Unmask 66.87 55.75 78.01 77.84 61.88 81.41
AugGPT 66.08 54.80 78.28 77.79 60.18 79.18
Grammar 66.20 55.75 76.96 76.52 57.80 81.03
Spelling 66.24 55.53 78.16 77.57 62.93 81.99
Chain 65.85 54.83 78.04 77.98 59.88 81.50
Hint 66.41 55.75 77.52 78.06 61.13 81.38
Taboo 64.27 56.13 77.72 77.36 61.55 81.30
w/o Aug 67.85 52.79 79.47 79.74 53.18 81.63
w/o Coreset 65.24 50.89 76.98 77.92 51.85 81.01
w/o Selective 66.05 50.23 78.76 79.00 53.22 80.85
w/o DPO|DS 66.74 55.51 79.71 79.64 61.06 82.09
w/o DPO 67.05 54.97 79.72 79.51 61.79 81.83
w/o Sampling 67.50 55.93 79.79 78.97 62.03 82.18
DoAug 68.14 56.25 79.83 79.93 63.91 82.20
Table 4: Training the GPT-2 and T5-large models on CoLA, MNLI, and RCT dataset.
(a) Average accuracy
(b) Distance
(c) Dispersion
Figure 5: Performance and diversity comparison between Llama and Qwen.
4.5 LLM Architectures Exploration
To exhibit the generalizability of our proposed methodology, we replace the LLM augmenter and downstream task model with other LLM architectures respectively. The results show that DoAug is agnostic to LLM architectures.
For the LLM augmenter, we replace the Llama-3.2-1B-Instruct model with the similar-sized Qwen2.5-1.5B-Instruct model Team (2024). As shown in Figure 5, datasets augmented by the Qwen model significantly outperform the original dataset and achieve comparable accuracy and diversity with those of the Llama model. Detailed results are given in Appendix F.
For the downstream task model, we replace BERT, an encoder-only model with the GPT-based and T5-based classification models. GPT is a decoder-only LLM and is especially adept at generating texts from a prompt. Based on an encoder-decoder transformer architecture, T5 is trained to perform all NLP tasks in a unified text-to-text format and is favorable in broad cases. Specifically, we use GPT-2 Radford et al. (2019) and T5-large Raffel et al. (2020) as the backbone of classification models, train these models on the MNLI, CoLA, and RCT datasets, and collect their performances. Experimental results in Table 4 show that DoAug  benefits both decoder-only models such as GPT and encoder-decoder models such as T5.
5 Conclusion
Diversity is an important factor in developing AI-ready, high-quality datasets but is often ignored in data augmentation methods. We propose a Diversity-oriented data Augmentation framework (DoAug) that trains an LLM paraphraser to enlarge and introduce diversity to textual datasets. The LLM paraphraser is fine-tuned to rewrite existing sentences in the original datasets, generating high-affinity samples and ensuring coherence of the dataset. We further construct a preference dataset and then fine-tune the LLM paraphraser with the DPO algorithm to encourage diversified generation. In this way, we maximize the diversity of the augmented dataset in our method. In extensive experiments, our proposed method exhibits a remarkable capability to boost dataset diversity, and the diversity gain significantly benefits the model’s learning performance of downstream tasks.
6 Limitations
This study has several limitations that should be acknowledged and addressed in future work.
Diversity Exploration: The evaluation of diversity lacks a universally accepted metric. In this study, we employed a subset of diversity-related evaluation methods, but other metrics, such as n-gram-based measures (e.g., assessing uniqueness or non-repetition), were not utilized. This limitation suggests that our assessment of diversity may not fully capture all aspects of the concept.
Augmentation Validation: Evaluating the correctness of generated data remains a challenging task. While both human evaluation and model-assisted evaluation are viable approaches, each comes with its own limitations. In this study, we relied on a sentiment-neutral large language model (LLM) for evaluation. However, this approach has inherent constraints, such as potential biases in the model and its inability to fully capture nuanced correctness in certain contexts.
Generation Factors: The quality and characteristics of generated samples are influenced by multiple factors, including the choice of prompts and the specific LLMs used. In this study, the prompts and LLMs were limited in number, and we did not exhaustively explore all possible configurations. This restriction may have impacted the diversity and quality of the generated samples.
Evaluation Benchmarks: Our evaluation was primarily focused on sentence classification tasks, and we did not extend our analysis to more general tasks, such as mathematical reasoning, instruction-following, creative writing, question-answering (QA), or chain-of-thought (CoT) reasoning. Additionally, we did not explore multimodality scenarios, which could provide a broader perspective on the applicability of our framework. Furthermore, the datasets used in this study were limited to English corpora, and we did not consider multilingual datasets, which could offer insights into cross-lingual or language-specific performance.
Potential Risks of Using LLM: Leveraging LLMs for data augmentation might suffer from demographic bias and factual inaccuracies. First, LLMs may amplify demographic biases from their training data. Second, generation hallucinations may produce plausible but factually incorrect content. When task models train on such flawed data, their reliability and accuracy degrade, especially in high-stakes domains like healthcare or finance. Mitigating these risks requires rigorous validation, bias-detection frameworks, and human oversight to ensure the generated datasets uphold fairness and factual integrity.
These limitations highlight areas for future work, including the adoption of more comprehensive diversity metrics, broader validation approaches, exploration of additional prompts and LLMs, and evaluation across diverse tasks, modalities, and languages.
Appendix A Dataset Specification
The details of the 12 NLP datasets used in our experiments, including the domain, application task, input scheme, and class number, are summarized in Table 5.
Appendix B Implementation Details
We use Llama-3.2-1B-Instruct with BF16 quantization as the LLM paraphraser. Llama-3.2 is one of the latest products of the Llama family. The Llama-3.2-1B model outperforms the Llama-3.2-3B and Llama-3.1-8B models on the rewriting task while requiring minimal memory and inference time [^2] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct. We fine-tune the Llama model with the LlamaFactory framework Zheng et al. (2024). In the SFT stage, the learning rate is $10^{-4}$. In the DPO stage, $\beta$ in the loss function is set to $0.1$ and the learning rate is $5^{-6}$. The LLM is trained for 3 epochs with the AdamW optimizer, a cosine scheduler, and a warm-up ratio of $0.1$ in each stage. The rank $r$ for LoRA fine-tuning is $8$. $\mathcal{D}_{\text{SFT}}$ and $\mathcal{D}_{\text{DPO}}$ contain 100,000 sentence pairs and 50,000 preference pairs, respectively. We use the embedding vector of the [CLS] token in the last layer of the $\text{BERT}_{\text{base}}$ model as the embedding space $\mathcal{E}$ for both preference dataset construction and dataset diversity evaluation. In the coreset selection step, the coreset ratio $r_{\text{augment}}:r_{\text{retain}}:r_{\text{prune}}$ is $1:1:1$. For each original sentence, we generate $K=5$ paraphrases and sample the most diversified output. For downstream task evaluation, we train the BERT model for 3 epochs. The model is updated with the AdamW optimizer. The learning rate is $5^{-5}$ with a linear scheduler and no warm-up. The LLM paraphraser and downstream task models are trained on two A100-40G GPUs with transformers 4.45.2, pytorch 2.5.1, and CUDA 12.4. It takes 32 and 36 minutes to train with SFT and DPO, respectively, and the LLM augmenter can paraphrase roughly 1 sentence per second.
Appendix C Baseline Methods Implementation
The error ratios for OCR and Keyboard are set to $0.15$. The ratios for its EDA’s four operations in our experiments are set to $0.1$, respectively. For BT, we use the en-zh and zh-en versions of the opus-mt model by Helsinki-NLP Tiedemann and Thottingal (2020) as the translator. The masking ratio for Unmask is set to $0.15$ and we sample the top-1 predictions of the $\text{BERT}_{\text{base}}$ model. Considering the cost of calling ChatGPT APIs, when implementing AugGPT, Grammar, Spelling, Chain, Hint, and Taboo, we use the open-source Llama-3.1-8B-Instruct model Dubey et al. (2024) as a replacement. Since it achieves competitive performance compared with the GPT 3.5 Turbo model (e.g. 69.4 v.s. 70.7 on MMLU 5-shot and 80.4 v.s. 69.9 on IFEval) Dubey et al. (2024), replacing GPT with Llama hardly compromises the effectiveness of the baseline methods.
Appendix D Augmentation Examples
We include some augmentation examples of DoAugand baseline methods in Table 6. We can observe that DoAugintroduces more details (have trouble moving) and some novel vocabularies (fondness, sophisticated).
Appendix E Parameter Sensitivity Study
Given the heavy time and computation cost of training LLMs, we follow most settings from existing research or default configurations from the library. Still, we conduct sensitivity studies on some key parameters unique to our method.
E.1 Number of Generated Sentences.
First, we study the effect of $K$, the number of total sentences generated by the LLM paraphraser when paraphrasing an original sentence. As Figure 6a shows, the best performances are achieved when $K$ is between $5$ to $8$. Intuitively, too small $K$ limits the possibility and diversity of generated sentences and therefore affects the dataset diversity and task performance; on the other hand, too large $K$ is likely to allow the LLM paraphraser go too far from the original semantics, breaking label preservability.
E.2 Coreset Ratio.
We then use the MNLI dataset to study the effect of coreset ratio for data pruning and data augmentation. As presented in Figure 6b and Figure 6c, both data pruning and data augmentation favor a moderate ratio, where data pruning ratio is embodied by $r_{\text{prune}}/(r_{\text{prune}}+r_{\text{retain}}+r_{\text{augment}})$, and data augmentation is embodied by $r_{\text{augment}}/(r_{\text{retain}}+r_{\text{augment}})$.
E.3 Coreset Method.
The choice of coreset methods is another factor that influences the final performance. So, we also investigate how the performance changes when augmentation is applied to different coresets. Figure 7 shows that different datasets favor different coreset methods, however, in most cases, the suboptimal choices of coreset still outperform data augmentation without targeting coresets, and augmentation performance does not significantly degenerate on most suboptimal coresets. This result demonstrates our coreset-focused selective data augmentation method can benefit from appropriate coresets but is robust against suboptimal coresets.
(a) Generated sentences
(b) Pruning ratio
(c) Augmentation ratio
Figure 6: Parameter sensitivity study on generated sentence count, coreset ratio for data pruning and augmentation on MNLI dataset.
(a) CoLA
(b) RCT
(c) SST-2
(d) SUBJ
Figure 7: Sensitivity study on the choice of coreset method.
Figure 8: Qwen performance on 12 datasets
Appendix F Full Results of LLM Architecture Exploration
Detailed results of performance on all 12 datasets and diversity in terms of 5 metrics are given in Figure 8 and Figure 9.
Ch.Pr. CoLA MNLI RCT SST-2 SUBJ Sympt. YELP
600 Original Samples 58.33 74.56 42.81 71.62 86.97 95.73 74.06 51.48
1.2 K Original Samples 71.12 77.61 61.63 78.96 88.56 95.77 92.45 56.48
DoAug 70.22 75.62 59.76 80.10 88.64 95.80 90.74 56.57
Table 7: Results on low-resource datasets (600), larger original subsets (1.2 K), and DoAug (600 Original + 600 Augmented Samples).
(a) Distance
(b) Dispersion
(c) Distance
(d) Average accuracy
(e) Dispersion
Figure 9: Diversity comparison between Llama and Qwen
Appendix G Alleviating the Low-resource Problem
In our main experiments, we artificially create low-resource conditions by sampling a subset from the original dataset. In Table 7, we compare the results of DoAug against model performance on large subsets, which are two times the size of that used in our main experiments. The comparison shows that DoAug can alleviate the low-resource problem and even performs better than larger subsets in some cases.